A system containing m = 5 kg of a liquid is stirred with a torque of T = 0.3 kgf-m
at a speed of X = 1000 RPM for t = 24 hr. The system meanwhile expands from
V1= 1m³ to V2 = 2m³ against a constant pressure of p = 4 kgf/cm². Determine
the net work done (W) in kJ.
1
Expert's answer
2020-10-02T07:08:04-0400

Pressure is given, "P = 4kgf/cm^2 = 4000 \times 9.8 \times 10000 N/m^2 = 39.2 \times 10^7 N/m^2"

Initial Volume, "V_i = 1m^3"

Final Volume, "V_f = 2 m^3"

Work done by against pressure, "W_p = P(V_f - V_i) = 39.2\times 10^7 (2-1) J =39.2\times 10^7 J"



Torque acting on the water, "\tau = 0.3 kgfm = 0.3 \times 1000 \times9.8 = 2940 Nm"

Angular speed , "\omega = 1000RPM = \frac{2 \pi \times1000}{60 } rad/s = 104.72" "rad\/s"

time. "t = 24 hrs = 24 \times 3600 s = 86400s"

Work done by torque, "W_{\tau} = \tau \times \theta =\tau \times \omega t= (2940 \times 104.72 \times 86400) = 2660.1 \times 10^7 J"

Total work done, "W = W_p + W_{\tau} = (39.2 + 2660.1)\times 10^7 J = 2699.3\times 10^7 J"